# Jackknife Empirical Likelihood Method for U Statistics Based on Multivariate Samples and its Applications

### v1

@@ 3 Simulation
We conduct simulation studies with different sample sizes and simulation is repeated 1,000 time. We use R package “empilik” for finding JEL based confidence intervals. For constructing confidence intervals based on kernel based method we use 100 bootstrap replications. Two bivariate probability models are considered for the simulation studies.
Marshall-Olkin bivariate exponential model (MOBVE($\lambda_{1},\lambda_{2},\lambda_{3}$)) : We consider the Marshall-Olkin bivariate exponential model MOBVE($\lambda_{1},\lambda_{2},\lambda_{3}$), introduced by [13]. The joint cumulative distribution function of this bivariate model is given by
@@
In Table 2, we explored various configurations for sample sizes $(n_{1},n_{2},n_{3})$ and the parameters $(\lambda_{x_{1}},\lambda_{x_{2}},\lambda_{x_{3}};\lambda_{y_{1}},\lambda_{y_{2}},\lambda_{y_{3}};\lambda_{z_{1}},\lambda_{z_{2}},\lambda_{z_{3}})$. We employed the JEL, normal approximation, and kernel-based methods to compute the CP and AL of confidence intervals for $\theta$. This comprehensive investigation aims to evaluate the performance and efficacy of these methods under various parameter settings within the MOBVE framework.
@@
For $-1\leq\theta\leq 1$, the FGM copula, or joint C.D.F., is given by
@@
$\qquad\qquad\qquad 0\leq F_{X_{1}}(x_{1}),\,F_{X_{2}}(x_{2})\leq 1$
@@
Table 1: Simulation time taken by various methods under two models
@@
Table 2: Coverage Probability (CP) and Average length of confidence intervals for the difference two VUS’s at 95% confidence level : Under MOBVE models
$(\lambda_{x_{1}},\lambda_{x_{2}},\lambda_{x_{3}};\lambda_{y_{1}},\lambda_{y_{2}},\lambda_{y_{3}};\lambda_{z_{1}},\lambda_{z_{2}},\lambda_{z_{3}})$ $(n_{1},n_{2},n_{3})\;\;\;$ JEL Normal Approx Kernel
CP (%) AL CP (%) AL CP (%) AL
(1,1,1;1,1,1;1,1,1) (10,10,10) 96.8 .501 97.8 0.384 97 0.392
(20,20,20) 96.5 .332 96.8 0.325 95.8 0.341
(30,30,30) 95.7 .229 97 0.257 96.6 0.264
(50,50,50) 95 .18 96.6 0.183 96 0.195
(70,70,70) 94.4 .126 94 0.124 93.5 0.129
(100,100,100) 95.8 .102 94.2 0.103 94 0.122
(1,2,0;1,1,0;2,1,0) (10,10,10) 97 .564 92.8 0.516 91.8 0.504
(20,20,20) 96.4 .35 93 0.361 92 0.341
(30,30,30) 95.8 .248 92.5 0.239 93 0.25
(50,50,50) 95.4 .187 94 0.187 92.5 0.169
(70,70,70) 95.2 .157 94 0.187 92.5 0.158
(100,100,100) 95.2 .131 94.5 0.130 94 0.135
(1/3,1/3,2/3;2/3,2/3,4/3;1,1,2) (10,10,10) 99.4 .263 99.6 0.234 97.5 0.21
(20,20,20) 97 .169 97.5 0.151 98.5 0.161
(30,30,30) 96 .112 97.5 0.121 98.25 0.125
(50,50,50) 95.6 .085 95.75 0.096 97.5 0.089
(70,70,70) 95.6 .067 94.5 0.069 96.5 0.071
(100,100,100) 95.2 .053 94.8 0.058 96 0.062
(3/5,3/5,2/5;6/5,6/5,4/5;9/5,9/5,6/5) (10,10,10) 99.2 .319 98.75 0.269 96.5 0.235
(20,20,20) 95.2 .2 97 0.197 97.25 0.211
(30,30,30) 97 .146 96.5 0.151 97 0.142
(50,50,50) 96.4 .103 96.5 0.098 97 0.106
(70,70,70) 96 .087 96 0.082 96.5 0.09
(100,100,100) 95.6 .066 96 0.069 96 0.065
(1/19,1/19,18/19;2/19,2/19,36/19;3/19,3/19,54/19) (10,10,10) 98 0.09 86 0.07 88 0.079
(20,20,20) 95.2 0.071 98 0.06 98 0.055
(30,30,30) 97 0.048 99 0.052 94 0.049
(50,50,50) 91 0.034 89 0.034 88.5 0.036
(70,70,70) 90.4 0.027 90 0.028 88 0.029
(100,100,100) 93.6 0.023 91 0.022 90 0.025
@@
Table 3: Coverage Probability (CP) and Average length of confidence intervals for the difference two VUS’s at 95% confidence level : Under Bivariate Pareto models
$(\lambda_{x_{1}},\lambda_{x_{2}},\alpha_{x_{1}},\alpha_{x_{2}},\theta_{x};\lambda_{y_{1}},\lambda_{y_{2}},\alpha_{y_{1}},\alpha_{y_{2}},\theta_{y};\lambda_{z_{1}},\lambda_{z_{2}},\alpha_{z_{1}},\alpha_{z_{2}},\theta_{z})$ $(n_{1},n_{2},n_{3})\;\;\;$ JEL Normal Approx Kernel
CP (%) AL CP (%) AL CP (%) AL
(1,1,1,1,0.5;1,1,1,1,0.5;1,1,1,1,-0.5) (10,10,10) 96 .493 96.3 0.451 98 0.485
(20,20,20) 97 .311 97 0.298 96.6 0.322
(30,30,30) 95 .274 94.3 0.262 94.1 0.281
(50,50,50) 96.8 .192 93.8 0.201 93.7 0.182
(70,70,70) 95.2 .156 94.8 0.153 94 0.159
(100,100,100) 95.4 .109 94.8 0.110 94.2 0.101
(1,2,1,2,0.5;2,1,2,1,0.5;2,2,1,1,-0.5) (10,10,10) 96 .557 91 0.549 89 0.514
(20,20,20) 98 .362 99 0.38 91 0.391
(30,30,30) 92 .268 92 0.306 91 0.321
(50,50,50) 98 .193 99 0.235 93 0.242
(70,70,70) 97 .198 94 0.189 92.5 0.191
(100,100,100) 96 .183 94 0.185 94 0.187
(5,1,1,5,0.2;5,1,2,2,0.5;1,5,5,1,0.9) (10,10,10) 97 .460 90 0.465 88 0.45
(20,20,20) 93 .345 92 0.325 91 0.34
(30,30,30) 94 .275 92 0.295 92 0.29
(50,50,50) 96 .174 97 0.17 98 0.183
(70,70,70) 96 .165 93.5 0.166 94 0.169
(100,100,100) 95.7 .151 94 0.150 94 0.153
(0.5,1,0.5,1,-0.5;1,1,0.5,0.5,-0.5;1,1,5,1,-0.2) (10,10,10) 94 .359 92 0.356 91 0.341
(20,20,20) 98 .212 93 0.201 92 0.204
(30,30,30) 97 .169 93 0.157 92 0.17
(50,50,50) 96 .127 93 0.13 92.5 0.129
(70,70,70) 95.8 .098 93.5 0.095 93 0.097
(100,100,100) 95.5 .076 94 0.075 93.5 0.076
(10,1,0.5,10,-0.9;1,1,15,0.2,-0.1;0.2,1,5,0.1,0.9) (10,10,10) 92 .530 90 0.521 89 0.532
(20,20,20) 98 .317 91 0.29 89 0.299
(30,30,30) 97 .262 91 0.267 91 0.276
(50,50,50) 96 .196 92 0.196 93 0.189
(70,70,70) 96 .163 98 0.165 99 0.17
(100,100,100) 95.8 .123 94 0.12 93.5 0.125
The following observations can be made from Tables 1-3:
2. In most cases, as the sample size grows, the coverage probabilities of confidence intervals generated by the three methods tend to align with the nominal level (95%). This is anticipated because larger sample sizes yield more detailed data insights.
3. From Tables 2-3, we can see that CI based on JEL method achieves closer to 95% coverage across various conditions compared to the other two methods. The average interval lengths are generally better for CIs based on the normal approximation and kernel-based methods for small sample sizes. However, the average lengths are almost same for moderate sample sizes.
4 Data Analysis
@@
In total, 883 results were included in this analysis, with sample sizes for the three individual classes being 222, 122, and 539, respectively. We used the JEL, Normal approximation, and Kernel-based methods, as described in Section 2.1, to construct confidence intervals for the difference between two VUSs. The JEL method took 3.4 hours, the Normal approximation method took 4.7 hours, and the Kernel-based method (with 1000 bootstraps) took 52 hours to obtain results for this real data analysis using R software on the previously specified laptop.
@@
Table 4: Confidence Intervals for $\theta$ at different confidence levels using different methods
@@
5 JEL for $k\;(>3)$ Multivariate Sample U-statistics
Until now, the results developed in this paper have been restricted to three-sample-based U-statistics. However, in practical applications, there are many cases where U-statistics are based on more than three samples ([10]). To provide an alternative non-parametric method for estimation and testing purposes, in this section, we develop the JEL method for more than three multivariate samples based U statistics. Consider the following independent random samples
$\left(\mathbb{X}_{1}^{(1)},\ldots,\mathbb{X}_{n_{1}}^{(1)}\right),\left(\mathbb{X}_{1}^{(2)},\ldots,\mathbb{X}_{n_{2}}^{(2)}\right),\dots,\left(\mathbb{X}_{1}^{(k)},\ldots,\mathbb{X}_{n_{k}}^{(k)}\right)$
from $k$ independent $q$-dimensional distribution functions $F_{1},F_{2},\dots,F_{k}$, respectively, where
$\mathbb{X}_{i}^{(1)}$ $=(X_{11}^{(1)},X_{12}^{(1)},\dots,X_{1q}^{(1)})^{T},\;i=1,2,...,n_{1},$
$\mathbb{X}_{i}^{(2)}$ $=(X_{11}^{(2)},X_{12}^{(2)},\dots,X_{1q}^{(2)})^{T},\;i=1,2,...,n_{2},$
$\vdots$
$\mathbb{X}_{i}^{(k)}$ $=(X_{11}^{(k)},X_{12}^{(k)},\dots,X_{1q}^{(k)})^{T},\;i=1,2,...,n_{k}.$
Let $\theta=E\left[h\left(\mathbb{X}_{i_{1,1}}^{(1)},\ldots,\mathbb{X}_{i_{1,m_{1}}}^{(1)},\mathbb{X}_{i_{2,1}}^{(2)},\ldots,\mathbb{X}_{i_{2,m_{2}}}^{(2)},\dots,\mathbb{X}_{i_{k,1}}^{(k)}\ldots,\mathbb{X}_{i_{k,m_{k}}}^{(k)}\right)\right]\in\Re$ be the parameter of interest. A $k$-sample U-statistics of degree $\left(m_{1},m_{2},\dots,m_{k}\right)$ with a symmetric kernel $h$ is given by
$U_{n_{1},n_{2},\dots,n_{k}}=$ $d^{*}\hskip-7.22743pt\mathop{\sum\sum\cdots\sum}_{\begin{subarray}{c}1\leq i_{1,1}<\ldots<i_{1,m_{1}}\leq n_{1}\\ 1\leq i_{2,1}<\ldots<i_{2,m_{2}}\leq n_{2}\\ \vdots\;\;\;\\ 1\leq i_{k,1}<\ldots<i_{k,m_{k}}\leq n_{k}\end{subarray}}\hskip-7.22743pth\left(\mathbb{X}_{i_{1,1}}^{(1)},\ldots,\mathbb{X}_{i_{1,m_{1}}}^{(1)},\mathbb{X}_{i_{2,1}}^{(2)},\ldots,\mathbb{X}_{i_{2,m_{2}}}^{(2)},\dots,\mathbb{X}_{i_{k,1}}^{(k)}\ldots,\mathbb{X}_{i_{k,m_{k}}}^{(k)}\right),$
where
$d^{*}=\frac{1}{{n_{1}\choose m_{1}}{n_{2}\choose m_{2}}\cdots{n_{k}\choose m_{k}}},$
and $\{i_{1,1},\ldots,i_{1,m_{1}}\}$, $\{i_{2,1},\ldots,i_{2,m_{2}}\},\dots,\{i_{k,1},\ldots,i_{k,m_{k}}\}$ indicates $m_{1},m_{2},\dots,m_{k}$ integers chosen without replacement from the sets $\{1,\ldots,n_{1}\},\{1,\ldots,n_{2}\},\dots,\{1,\ldots,n_{k}\}$, respectively.
Let $n=\sum_{l=1}^{k}n_{l}$ and $m=\sum_{l=1}^{k}m_{l}$. Define $\mathbb{W}_{l},$ as
$\mathbb{W}_{l}=\begin{cases}\mathbb{X}_{l}^{(1)},&\text{ if }l=1,\dots,n_{1}\\ \mathbb{X}_{l-n_{1}}^{(2)},&\text{ if }l=n_{1}+1,\dots,n_{1}+n_{2}\\ \vdots\\ \mathbb{X}_{l-\sum_{i=1}^{k-1}n_{i}}^{(k)},&\text{ if }l=\sum_{i=1}^{k-1}n_{i}+1,\dots,n\end{cases}$
Therefore,
$(\mathbb{W}_{1},\mathbb{W}_{2},\dots,\mathbb{W}_{n})=\left(\mathbb{X}_{1}^{(1)},\ldots,\mathbb{X}_{n_{1}}^{(1)},\mathbb{X}_{1}^{(2)},\ldots,\mathbb{X}_{n_{2}}^{(2)},\dots,\mathbb{X}_{1}^{(k)},\ldots,\mathbb{X}_{n_{k}}^{(k)}\right).$
In same lines as the three sample case (in Section 2), we define the jackknife pseudo-values for combined single sample $(\mathbb{W}_{1},\mathbb{W}_{2},\dots,\mathbb{W}_{n})$, $V_{l},$ as
$V_{l}=nU_{n}-(n-1)U_{n-1}^{(-l)},\;\;\forall\;\;l=1,2,\dots,n,$ (17)
where $U_{n}$ is the U statistics based on single sample $(\mathbb{W}_{1},\mathbb{W}_{2},\dots,\mathbb{W}_{n})$ with degree $m$ and $U_{n-1}^{(-l)}$ is the U statistics based on sample $(\mathbb{W}_{1},\mathbb{W}_{2},\dots,\mathbb{W}_{n})$, after deleting observation $\mathbb{W}_{l}$. Note that $U_{n}=U_{n_{1},n_{2},\dots,n_{k}}$. It is easy to verify that $U_{n}=\frac{1}{n}\sum_{i=1}^{n}V_{i},$ where $V_{i}$ ’s, as defined by (17), are asymptotically independent random variables ([21]).
The values of $EV_{l}$ are as follows
$EV_{l}=\begin{cases}\theta\left(\frac{n}{n-m}\right)\left[n-m-\frac{(n-1)(n_{1}-m_{1})}{n_{1}}\right],&\text{ if }\;l=1,2,\ldots,n_{1}\\ ~{}\\ \theta\left(\frac{n}{n-m}\right)\left[n-m-\frac{(n-1)(n_{2}-m_{2})}{n_{2}}\right],&\text{ if }\;l=n_{1}+1,n_{1}+2,\ldots,n_{1}+n_{2}\\ \qquad\qquad\qquad\cdot&\\ \qquad\qquad\qquad\cdot&\\ \qquad\qquad\qquad\cdot&\\ \theta\left(\frac{n}{n-m}\right)\left[n-m-\frac{(n-1)(n_{k}-m_{k})}{n_{k}}\right],&\text{ if }\;l=\sum_{i=1}^{k-1}n_{i}+1,\ldots,n\end{cases},$
It is easy to verify that $\frac{1}{n}\sum_{l=1}^{n}EV_{l}=\theta$ and, in particular, if $m_{1}=m_{2}=\dots=m_{k}$ and $n_{1}=n_{2}=\dots=n_{k}$, we get $EV_{l}=\theta$, $l=1,2,\ldots,n$.
Let $\mathbf{p}=\left(p_{1},\ldots,p_{n}\right)$ be a probability vector assigning probability $p_{i}$ to each $V_{i}$. We know that $\prod_{i=1}^{n}p_{i}$ subject to $\sum_{i=1}^{n}p_{i}=1$ attain its maximum at $\frac{1}{n^{n}}$ at $p_{i}=\frac{1}{n}$. Hence the jackknife empirical likelihood ratio for $\theta$ is given by
$R(\theta)=\max\left\{\prod_{i=1}^{n}\left(np_{i}\right);\sum_{i=1}^{n}p_{i}=1,\sum_{i=1}^{n}p_{i}\left(V_{i}-EV_{i}\right)=0\right\}.$
Using Lagrange multipliers method, we obtain the jackknife empirical log-likelihood ratio as
$\log R(\theta)=-\sum_{i=1}^{n}\log\left(1+\lambda\left(V_{i}-EV_{i}\right)\right).$ (18)
Here also we can construct a confidence interval based on Wilks’ theorem. To state the result, we define
$\sigma_{i}^{2}$ $=Var\left[E\left(h\left(\mathbb{X}_{1}^{(1)},\ldots,\mathbb{X}_{m_{1}}^{(1)},\mathbb{X}_{1}^{(2)},\ldots,\mathbb{X}_{m_{2}}^{(2)},\dots,\mathbb{X}_{1}^{(k)},\ldots,\mathbb{X}_{m_{k}}^{(k)}\right)\mid\mathbb{X}_{1}^{(i)}\right)\right],\;\;i=1,2,\dots,k.$
We state the theorem which provides the asymptotic distribution of the jackknife empirical log-likelihood ratio. The proof of the theorem is on the same lines as of Theorem 2.2, and hence omitted.
Theorem 5.1.
Under the following conditions:
(a) $E\left(h^{2}\left(\mathbb{X}_{1}^{(1)},\ldots,\mathbb{X}_{m_{1}}^{(1)},\mathbb{X}_{1}^{(2)},\ldots,\mathbb{X}_{m_{2}}^{(2)},\dots,\mathbb{X}_{1}^{(k)},\ldots,\mathbb{X}_{m_{k}}^{(k)}\right)\right)<\infty$ and $\sigma_{i}^{2}>0,\;\forall\;i=1,2,\dots,k$;
(b) $0<\lim\inf\frac{n_{i-1}}{n_{i}}\leq\lim\sup\frac{n_{i-1}}{n_{i}}<\infty,\;\forall\;i=2,3,\dots,k$,
as $n\rightarrow\infty,-2\log R(\theta)$ converges in distribution to a $\chi^{2}$ random variable with one degree of freedom.
Using Theorem 5.1, the JEL based confidence interval for $\theta$ at $(1-\alpha)$ confidence level is given by
$CI=\{\theta\;|\;-2\log R(\theta)\leq\chi^{2}_{1,1-\alpha}\}.$
Also, the normal approximation based a two-sided confidence interval for $\theta$ at $(1-\alpha)$ confidence level is given by
$\left(U_{n_{1},n_{2},\dots,n_{k}}-z_{\alpha/2}\hat{\tau},U_{n_{1},n_{2},\dots,n_{k}}+z_{\alpha/2}\hat{\tau}\right),$
where $\hat{\tau}^{2}$ is estimator of asymptotic variance of $U_{n_{1},n_{2},\dots,n_{k}}$ that can be obtained in a similar lines as of $\hat{\sigma}^{2}$ in (2).
5.1 Application : Hypervolume Under the ROC Manifold
ROC curves are generalized to ROC manifolds for multi-category classification. The analogous extension of AUC is termed the Hypervolume Under the ROC Manifold (HUM). [10] developed a method for estimating HUM in unordered classifications by using a probabilistic interpretation. They also applied HUM as a criterion for model selection in microarray studies. For more details on HUM, refer to Chapter 3 of [25]. In this section, to illustrate the application of the preceding section, we consider constructing confidence intervals for the difference of two dependent HUMs in k-category classification.
Let
$\left(\mathbb{X}_{1}^{(1)},\ldots,\mathbb{X}_{n_{1}}^{(1)}\right),\left(\mathbb{X}_{1}^{(2)},\ldots,\mathbb{X}_{n_{2}}^{(2)}\right),\dots,\left(\mathbb{X}_{1}^{(k)},\ldots,\mathbb{X}_{n_{k}}^{(k)}\right)$
be $k$ random samples having $k$ independent joint distribution functions $F_{1},F_{2},\dots,F_{k}$, respectively, where
$\mathbb{X}_{i}^{(j)}$ $=(X_{11}^{(j)},X_{12}^{(j)})^{T},\;i=1,2,...,n_{j},\;\;j=1,2,\dots,k.$
Difference of two $k$-category classification dependent HUM is defined as
$\theta=P\left(X_{11}^{(1)}<X_{11}^{(2)}<\cdots<X_{11}^{(k)}\right)-P\left(X_{12}^{(1)}<X_{12}^{(2)}<\cdots<X_{12}^{(k)}\right).$
Therefore, we have
$\theta=E\left[h^{HUM}\left(\mathbb{X}_{1}^{(1)},\mathbb{X}_{1}^{(2)},\dots,\mathbb{X}_{1}^{(k)}\right)\right],$
where
$h^{HUM}\left(\mathbb{X}_{1}^{(1)},\mathbb{X}_{1}^{(2)},\dots,\mathbb{X}_{1}^{(k)}\right)=I_{\left(X_{11}^{(1)}<X_{11}^{(2)}<\cdots<X_{11}^{(k)}\right)}-I_{\left(X_{12}^{(1)}<X_{12}^{(2)}<\cdots<X_{12}^{(k)}\right)}.$
In this case, we have $q=2$ and $m_{1}=m_{2}=\cdots=m_{k}=1,$ and therefore, a $k$-sample U-statistic of degree $\left(1,1,\dots,1\right)$ with a symmetric kernel $h^{HUM}$ is given by
[…]
Using $U_{n}^{HUM}$, we can find jackknife pseudo values $V_{l},\;\forall\;l=1,2,\dots,n,$. Then the log JEL ratio $\log R(\theta)$ can be evaluated using the equation (18). In view of Theorem 5.1, the JEL based $(1-\alpha)$ level confidence interval for $\theta$ is given by
$CI=\{\theta\;|\;-2\log R(\theta)\leq\chi^{2}_{1,1-\alpha}\}.$
And the normal approximation based a two-sided confidence interval for $\theta$ at $(1-\alpha)$ confidence level is given by
$\left(U_{n}^{HUM}-z_{\alpha/2}\hat{\tau},U_{n}^{HUM}+z_{\alpha/2}\hat{\tau}\right).$
6 Concluding remarks
In this paper, we have described the JEL method for $k\,(\geq 3)$ multivariate samples U-statistics. This study opens the door for a wider range of applications in multivariate scenarios. By formalizing the JEL approach for $q$-dimensional $k$-sample-based U-statistics, we have established the asymptotic distribution of the JEL ratio and successfully constructed confidence intervals for unknown parameters.
To demonstrate the applications of our findings, we constructed confidence intervals for differences in two volumes under ROC surfaces (VUSs) and differences in two hypervolumes under the ROC manifolds (HUMs). These applications highlight the practical utility of the JEL method in addressing complex statistical problems. To evaluate the performance of the JEL method in comparison with other non-parametric approaches, we conducted extensive simulation studies under various models, including the Marshall-Olkin bivariate exponential and Bivariate Pareto distributions. Our results showed the superior performance of the JEL method in terms of coverage probability and computational complexity when compared to the normal approximation using jackknife pseudo-values based variance estimator and kernel-based methods.
Additionally, we conducted a real data analysis using an Alzheimer’s disease dataset to illustrate the practical benefits of the JEL method in real-world scenarios. Overall, this study highlights the superior performance, versatility, lower computational cost and practical benefits of the JEL method, recommending its use in statistical practices that require reliable nonparametric inference for complex data structures.
Furthermore, the results of this paper facilitate hypothesis testing using the JEL method, based on the established asymptotic distribution of the JEL ratio. Previous studies, including [23], [12], and Chapter 8 of [3], have explored testing problems based on U-statistics. The JEL method we developed in this paper can be effectively applied to these types of problems.

### v2

@@ 4. Simulation
We conduct simulation studies with different sample sizes and simulation is repeated 1,000 time. We use R package “empilik” for finding JEL based confidence intervals. For constructing confidence intervals based on kernel based method we use 1000 bootstrap replications. Two bivariate probability models are considered for the simulation studies.
Marshall-Olkin bivariate exponential model (MOBVE($\lambda_{1},\lambda_{2},\lambda_{3}$)) : We consider the Marshall-Olkin bivariate exponential model MOBVE($\lambda_{1},\lambda_{2},\lambda_{3}$), introduced by marshall1967multivariate. The joint cumulative distribution function of this bivariate model is given by
@@
In Table 2, we explored various configurations for sample sizes $(n_{1},n_{2},n_{3})$ and the parameters $(\lambda_{x_{1}},\lambda_{x_{2}},\lambda_{x_{3}};\lambda_{y_{1}},\lambda_{y_{2}},\lambda_{y_{3}};\lambda_{z_{1}},\lambda_{z_{2}},\lambda_{z_{3}})$. This comprehensive investigation aims to evaluate the performance and efficacy of these three methods under various parameter settings within the MOBVE framework.
@@
For $1\leq\theta\leq 1$, the FGM copula, or joint C.D.F., is given by
@@
$\qquad\qquad\qquad 0\leq F_{X_{1}}(x_{1}),\,F_{X_{2}}(x_{2})\leq 1.$
@@
The following observations can be made from Tables 1- 3:
2. In most cases, as the sample size increases, the coverage probabilities of the confidence intervals generated by the three methods tend to align with the nominal level (95%). This is anticipated because larger sample sizes yield more detailed data insights.
3. From Tables 2-3, we can see that CI based on JEL method achieves closer to 95% coverage across various choices of parameters compared to the other two methods. The average interval lengths are generally better for CIs based on the normal approximation and kernel-based methods for small sample sizes. However, the average lengths are almost same for moderate sample sizes.
Note that an2018jackknife also conducted simulation studies for the confidence intervals defined in (3.1) and (3.2). However, they did not consider the bivariate Pareto model and did not compare these methods with kernel-based confidence intervals.
Table 1. Simulation time taken by various methods under two models
@@
Table 2. Coverage probability (CP) and average length (AL) of confidence intervals for the difference two VUS’s at 95% confidence level: Under MOBVE models
$(\lambda_{x_{1}},\lambda_{x_{2}},\lambda_{x_{3}};\lambda_{y_{1}},\lambda_{y_{2}},\lambda_{y_{3}};\lambda_{z_{1}},\lambda_{z_{2}},\lambda_{z_{3}})$ $(n_{1},n_{2},n_{3})\;\;\;$ JEL Normal Approx Kernel CP (%) AL CP (%) AL CP (%) AL (1,1,1;1,1,1;1,1,1) (10,10,10) 96.8 .501 97.8 0.384 97 0.392 (20,20,20) 96.5 .332 96.8 0.325 95.8 0.341 (30,30,30) 95.7 .229 97 0.257 96.6 0.264 (50,50,50) 95 .18 96.6 0.183 96 0.195 (70,70,70) 94.4 .126 94 0.124 93.5 0.129 (100,100,100) 95.8 .102 94.2 0.103 94 0.122 (1,2,0;1,1,0;2,1,0) (10,10,10) 97 .564 92.8 0.516 91.8 0.504 (20,20,20) 96.4 .35 93 0.361 92 0.341 (30,30,30) 95.8 .248 92.5 0.239 93 0.25 (50,50,50) 95.4 .187 94 0.187 92.5 0.169 (70,70,70) 95.2 .157 94 0.187 92.5 0.158 (100,100,100) 95.2 .131 94.5 0.130 94 0.135 (1/3,1/3,2/3;2/3,2/3,4/3;1,1,2) (10,10,10) 99.4 .263 99.6 0.234 97.5 0.21 (20,20,20) 97 .169 97.5 0.151 98.5 0.161 (30,30,30) 96 .112 97.5 0.121 98.25 0.125 (50,50,50) 95.6 .085 95.75 0.096 97.5 0.089 (70,70,70) 95.6 .067 94.5 0.069 96.5 0.071 (100,100,100) 95.2 .053 94.8 0.058 96 0.062 (3/5,3/5,2/5;6/5,6/5,4/5;9/5,9/5,6/5) (10,10,10) 99.2 .319 98.75 0.269 96.5 0.235 (20,20,20) 95.2 .2 97 0.197 97.25 0.211 (30,30,30) 97 .146 96.5 0.151 97 0.142 (50,50,50) 96.4 .103 96.5 0.098 97 0.106 (70,70,70) 96 .087 96 0.082 96.5 0.09 (100,100,100) 95.6 .066 96 0.069 96 0.065 (1/19,1/19,18/19;2/19,2/19,36/19;3/19,3/19,54/19) (10,10,10) 98 0.09 86 0.07 88 0.079 (20,20,20) 95.2 0.071 98 0.06 98 0.055 (30,30,30) 97 0.048 99 0.052 94 0.049 (50,50,50) 91 0.034 89 0.034 88.5 0.036 (70,70,70) 90.4 0.027 90 0.028 88 0.029 (100,100,100) 93.6 0.023 91 0.022 90 0.025
@@
Table 3. Coverage probability (CP) and average length (AL) of confidence intervals for the difference two VUS’s at 95% confidence level: Under Bivariate Pareto models
$(\lambda_{x_{1}},\lambda_{x_{2}},\alpha_{x_{1}},\alpha_{x_{2}},\theta_{x};\lambda_{y_{1}},\lambda_{y_{2}},\alpha_{y_{1}},\alpha_{y_{2}},\theta_{y};\lambda_{z_{1}},\lambda_{z_{2}},\alpha_{z_{1}},\alpha_{z_{2}},\theta_{z})$ $(n_{1},n_{2},n_{3})\;\;\;$ JEL Normal Approx Kernel CP (%) AL CP (%) AL CP (%) AL (1,1,1,1,0.5;1,1,1,1,0.5;1,1,1,1,-0.5) (10,10,10) 96 .493 96.3 0.451 98 0.485 (20,20,20) 97 .311 97 0.298 96.6 0.322 (30,30,30) 95 .274 94.3 0.262 94.1 0.281 (50,50,50) 96.8 .192 93.8 0.201 93.7 0.182 (70,70,70) 95.2 .156 94.8 0.153 94 0.159 (100,100,100) 95.4 .109 94.8 0.110 94.2 0.101 (1,2,1,2,0.5;2,1,2,1,0.5;2,2,1,1,-0.5) (10,10,10) 96 .557 91 0.549 89 0.514 (20,20,20) 98 .362 99 0.38 91 0.391 (30,30,30) 92 .268 92 0.306 91 0.321 (50,50,50) 98 .193 99 0.235 93 0.242 (70,70,70) 97 .198 94 0.189 92.5 0.191 (100,100,100) 96 .183 94 0.185 94 0.187 (5,1,1,5,0.2;5,1,2,2,0.5;1,5,5,1,0.9) (10,10,10) 97 .460 90 0.465 88 0.45 (20,20,20) 93 .345 92 0.325 91 0.34 (30,30,30) 94 .275 92 0.295 92 0.29 (50,50,50) 96 .174 97 0.17 98 0.183 (70,70,70) 96 .165 93.5 0.166 94 0.169 (100,100,100) 95.7 .151 94 0.150 94 0.153 (0.5,1,0.5,1,-0.5;1,1,0.5,0.5,-0.5;1,1,5,1,-0.2) (10,10,10) 94 .359 92 0.356 91 0.341 (20,20,20) 98 .212 93 0.201 92 0.204 (30,30,30) 97 .169 93 0.157 92 0.17 (50,50,50) 96 .127 93 0.13 92.5 0.129 (70,70,70) 95.8 .098 93.5 0.095 93 0.097 (100,100,100) 95.5 .076 94 0.075 93.5 0.076 (10,1,0.5,10,-0.9;1,1,15,0.2,-0.1;0.2,1,5,0.1,0.9) (10,10,10) 92 .530 90 0.521 89 0.532 (20,20,20) 98 .317 91 0.29 89 0.299 (30,30,30) 97 .262 91 0.267 91 0.276 (50,50,50) 96 .196 92 0.196 93 0.189 (70,70,70) 96 .163 98 0.165 99 0.17 (100,100,100) 95.8 .123 94 0.12 93.5 0.125
5. Data analysis
@@
In total, 883 results were included in this analysis, with sample sizes for the three individual classes being 222, 122, and 539, respectively. We used the JEL, Normal approximation, and Kernel-based methods, as described in Section 3, to construct confidence intervals for the difference between two VUSs. The JEL method took 3.4 hours, the Normal approximation method took 4.7 hours, and the Kernel-based method (with 1000 bootstraps) took 52 hours to obtain results for this real data analysis using R software on the previously specified laptop.
@@
Note that, for the application, we consider a recent dataset where the biomarkers and disease stages are entirely distinct from those used in the real-data analysis in an2018jackknife.
Table 4. Confidence Intervals for $\theta$ at different confidence levels using different methods
@@
6. Concluding remarks
In this paper, we studied U-statistics based on three multivariate samples and developed a jackknife empirical likelihood (JEL) methodology for such statistics. This would extend the applicability of JEL methods to three sample multivariate problems.
We constructed confidence intervals for differences in two dependent VUSs using three nonparametric procedures. We conducted an extensive simulation study under the Marshall-Olkin bivariate exponential and Bivariate Pareto distributions. Our results showed the superior performance of the JEL method in terms of coverage probability and computational complexity when compared to the normal approximation using jackknife pseudo-values based variance estimator and kernel-based methods.
Additionally, we conducted a real data analysis using a recent Alzheimer’s disease dataset to illustrate the practical benefits of the JEL method in real-world scenarios. This study highlights the superior performance, versatility, lower computational cost and practical benefits of the JEL method over its competitors recommending its use in statistical practices that require reliable nonparametric inference for complex data structures.
Beyond diagnostic testing, the proposed framework is also relevant in other scientific domains. For example, an ecologist could be interested in comparison of multiple species in three habitats: protected forest, logged forest and agricultural land. He would be interested in testing the hypothesis that the distribution of multiple species in these three habitats are identical. rizzo2016energy have discussed energy based statistics for this testing problem. Many of them can be expressed as U-statistics. Asymptotic distributions of the test statistics (either U-statistics or JEL ratio) will follow from the results in Section 2.
We are working on $q$-dimensional $k$-sample U-statistics and the asymptotic distribution of the associated JEL ratio. One possible application of the theory developed would be construction of confidence intervals for hypervolume under the ROC manifold which is a generalization of ROC curves to multi-category classification. The results will be presented elsewhere.